# Linzhan Mou5, Linqi Song4, Zhenan Sun [^2] Corresponding authors. 1,3, Ying Wei22footnotemark: 2 6 1

### v2

@@ Abstract
Quantization of large language models (LLMs) faces significant challenges, particularly due to the presence of outlier activations that impede efficient low-bit representation. Traditional approaches predominantly address Normal Outliers, which are activations across all tokens with relatively large magnitudes. However, these methods struggle with smoothing Massive Outliers that display significantly larger values, which leads to significant performance degradation in low-bit quantization. In this paper, we introduce DuQuant, a novel approach that utilizes rotation and permutation transformations to more effectively mitigate both massive and normal outliers. First, DuQuant starts by constructing rotation matrices, using specific outlier dimensions as prior knowledge, to redistribute outliers to adjacent channels by block-wise rotation. Second, We further employ a zigzag permutation to balance the distribution of outliers across blocks, thereby reducing block-wise variance. A subsequent rotation further smooths the activation landscape, enhancing model performance. DuQuant simplifies the quantization process and excels in managing outliers, outperforming the state-of-the-art baselines across various sizes and types of LLMs on multiple tasks, even with 4-bit weight-activation quantization. Our code is available at https://github.com/Hsu1023/DuQuant.
@@ 1 Introduction
Large language models (LLMs) [45, 7, 57, 18] have demonstrated exceptional performance across a wide range of natural language processing tasks. However, their billions of parameters present considerable deployment challenges on resource-constrained edge devices, particularly in terms of memory usage and inference speed [21, 16]. In response to these challenges, network quantization methods [19, 23] have been extensively explored to minimize memory usage by converting floating-point parameters into low-bit formats [17, 30, 8], and to expedite inference by quantizing both activations and weights for accelerating the matrix multiplication process [52, 31, 61].
@@
Extensive evaluations demonstrate that our DuQuant approach significantly outperforms existing 4-bit weight-activation quantization baselines across various benchmarks. Notably, DuQuant achieves a 5% improvement in Commonsense QA tasks across all LLaMA model sizes and a 10% increase in zero-shot MMLU benchmarks for the Vicuna-v1.5-13B. Moreover, in practical applications with the LLaMA2-7B model, DuQuant not only accelerates prefilling phase by up to 2.08$\times$ but also reduces memory usage by 3.20$\times$, with minimal impact on performance: only a 0.61 increase in perplexity and a 2.71% drop in accuracy compared to the FP16 model. These results highlight the effectiveness of DuQuant in enhancing both the efficiency and capacity of quantized LLMs.
@@ 2 Motivation
Previous works [12, 60, 30] have highlighted the challenge posed by activation outliers in LLMs for model compression. These outlier features consistently manifest large values across specific feature dimensions and are present in all token sequences [52], which we refer to as Normal Outliers. Recently, a distinct type of outlier [43, 32], termed Massive Outliers, also known as attention sinks [53], has been observed in LLMs. The primary distinctions between normal and massive outliers are: 1) Normal outliers persist across all token sequences, whereas massive outliers are confined to a limited number of tokens. 2) Massive outliers exhibit significantly larger magnitudes, often surpassing 100 and being approximately 1000 times greater than the median of other activations [43]. In our study, we delve deeper into the impact of these two distinct types of outliers on quantization.
@@
In contrast to previous studies [43, 32] that observe massive outliers at the output of Transformer blocks, we first discover that these extremely large activations exist at the input of the down-projection layer within the FFN module. As depicted in Figure 1, the input of the down-projection layer in the LLaMA2-7B model Layer 1 contains a single activation of significant magnitude (approximately 1200). This activation is isolated to one token and therefore classified as one of massive activations. This phenomenon is consistently observed across different layers and sizes of models, as illustrated in Appendix H.
@@
Although previous studies [52, 42, 36, 1] have proposed various approaches to eliminate outlier features, they still face challenges in effectively managing massive outliers. SmoothQuant [52], for instance, attempts to shift the quantization difficulty from activations to weights by dividing the activation by a per-channel smoothing factor and multiplying it to the weight matrix. Nevertheless, we observe that this transfer at the input of the down-projection layer can cause the weights of the down-projection to display noticeable outliers, as demonstrated in Figure 1 . This issue arises because massive outliers cause the smoothing factor to become significantly large. Moreover, extremely large outliers can lead optimization-based methods to encounter problems with loss explosion. Both OmniQuant [42] and AffineQuant [36] have had to exclude their learnable parameters for the down projection layer due to unstable gradients. Given the poor accuracy observed with 4-bit quantization, QUIK [1] opts to use INT8 quantization for the down projection layer and Atom [61] applies INT8 quantization for 128 outlier channels. Consequently, massive outliers introduce new challenges to the quantization process that existing methods cannot fully address. This observation has motivated us to develop rotation and permutation transformation, which effectively handles both massive and normal outliers and achieves state-of-the-art performance.
@@ 3.2 The proposed DuQuant Method
To address the Normal Outliers issue stated in Section 2, current quantization methods, such as SmoothQuant [52] and OmniQuant [52], usually adopt the smooth technique. Concretely, it involves the utilization of a per-channel smoothing diagonal matrix, denoted as $\mathbf{\Lambda}$, to scale the input activation and weight matrix. The adjustment allows us to rewrite the original linear layer as $\mathbf{Y}=\mathbf{X}\cdot\mathbf{W}=(\mathbf{X}\cdot\mathbf{\Lambda})(\mathbf{\Lambda}^{-1}\cdot\mathbf{W}).$ The diagonal element $\mathbf{\Lambda}_{j}$ within $\mathbf{\Lambda}$ is computed as $\mathbf{\Lambda}_{j}\!=\text{max}(|\mathbf{X}_{j}|)^{\alpha}/\text{max}(|\mathbf{W}_{j}|)^{1-\alpha}$, where $\alpha$ is a hyper-parameter representing the migration strength. However, despite the ability of this smoothing technique to shift the quantization challenge from activations to weights, it still faces difficulties in effectively managing Massive Outliers, as depicted in Figure 1. This challenge stems from the extremely large massive outliers inducing large scaling factors $\mathbf{\Lambda}_{j}$, which in turn introduce new outliers in the weight matrix and result in significant performance declines in low-bit quantization scenarios, such as INT4 .
@@
    Let $N$ denote the greedy search steps, then the approximated rotation matrix $\hat{\mathbf{R}}=\mathbf{R}^{1}\mathbf{R}^{2}\cdots\mathbf{R}^{n}$, where $n=\arg\min_{k\in[1:N]}\left(\max_{i,j}|(\mathbf{X}\mathbf{R}^{1}\cdots\mathbf{R}^{k})_{ij}|\right)$. Each $\mathbf{R}^{i}$ is constructed according to Eqn. (2) and the identified feature dimension $d^{(i)}$.
@@
where the notation $\mathbf{P}$ denote the orthogonal permutation matrix learned via the zigzag manner, the $\hat{\mathbf{R}}_{(1)}$ and $\hat{\mathbf{R}}_{(2)}$ represent the first and second block-diagonal rotation matrix, respectively.
Remark 1. It is worth noting that the proposed RAP method can simultaneously smooth the weight matrix. While the commonly adopted smooth technique is effective, it can cause the weight matrix of the down-projection layer to exhibit pronounced outliers, leading to performance degradation. However, in the proposed RAP method, the rotation transformation we designed is applied to not only the activation input but also the weight matrix. As a result, the outliers induced by the smooth technique can be mitigated through our approximated rotation matrix $\hat{\mathbf{R}}$, yielding a smoother, more quantization-friendly weight matrix. Moreover, this approach eliminates the reliance on complex weight quantization techniques, such as GPTQ [17] used in Atom [61].
@@ 4 Experiment
We apply our DuQuant on pre-trained LLMs: LLaMA (7B-65B) [45], LLaMA2 (7B-70B) [46], LLaMA3 (8B, 70B) and instruction-tuned LLMs: Vicuna-v1.5 (7B-13B) [9]. We evaluate quantized LLaMA models on language generation tasks and commonsense QA tasks. Specifically, we assess the perplexity on WikiText2 [37] and C4 [39] datasets, as well as the zero-shot accuracy on PIQA [6], ARC [11], BoolQ [10], HellaSwag [56], and WinoGrande [40] datasets. Moreover, we evaluate quantized Vicuna models on MMLU [20] and MT-Bench [62] benchmarks.
@@
In line with prior studies [31, 42], we apply per-token activation quantization and per-channel weight quantization. Given that W8A8 quantization has been established as lossless in precision by SmoothQuant [52], our primary evaluation in this paper focuses on 4-bit and 6-bit quantization for weights and activations. As for details, we quantize all intermediate activations, excluding the SoftMax output. Moreover, we have developed two types of quantized models, denoted as DuQuant and DuQuant+LWC . For DuQuant, we employ round-to-nearest quantization, using a clipping ratio of 0.9 for activations and 0.8 for weights. To improve weight matrix quantization, DuQuant+LWC integrates the learnable weight clipping (LWC) technique from OmniQuant. Concretely, LWC adjusts weights by training parameters $\gamma,\beta\in[0,1]$ to compute step size $\Delta=\frac{\gamma\max(\mathbf{X})-\beta\min(\mathbf{X})}{2^{b}-1}$ in Eqn. (1). Notably, the smoothing diagonal matrix and the learned weight clipping factor can be integrated into the quantized weights, introducing no additional computational or memory costs. More details and hyperparameters are left in Appendix C.
@@
Table 2: Zero-shot QA ($\uparrow$) results of LLaMA1 models under 4-bit weight-activation quantization. The results for LLaMA2 models and W6A6 quantization can be found in Table D1  D5, and  D6.
@@ 4.1 Main Results
We conduct a comprehensive comparison of our DuQuant with several SOTA baselines on LLaMA1 and LLaMA2 models. Results for W4A4 quantization are presented in this Section, while results for W6A6 quantization are provided in Appendix D. Table 1 indicates that our DuQuant quantized models notably outperform other baselines on both the WikiText2 and C4 datasets. Notably, LWC further enhances model capacity, with our DuQuant+LWC achieving comparable performance with FP16 models. Table 2 and Table D1 showcase the zero-shot accuracy of W4A4 quantization on Commonsense QA tasks, where DuQuant significantly improves the average accuracy. Our method surpasses QLLM by +9%, and Atom by +5% for all model sizes. These results demonstrate the superiority of our rotation and permutation transformation, which establishes new SOTA performance by effectively eliminating outlier features.
@@ 4.2 Ablation Study
Our DuQuant includes three main components: smoothing operation, initial rotation, and permutation with a second rotation. We combine these components to quantize the LLaMA2-13B model and evaluate their effects on language generation tasks. Results in Table 5 show that the smoothing operation plays a basic role in our DuQuant by shifting activation outliers to weight. The initial rotation significantly enhances model performance, yielding competitive PPL results. Finally, permutation combined with a second rotation further enhances the quantized model.
@@
In light of the recent introduction of Hadamard rotations by QuaRot[2] to eliminate outlier features, we have undertaken a detailed analysis to highlight the key differences between our DuQuant and QuaRot. To ensure a balanced evaluation, we have re-implemented QuaRot in accordance with our quantization settings. The results demonstrate that 1) the rotation matrix constructed by DuQuant outperforms QuaRot’s approach of simply selecting a randomly initialized Hadamard matrix. As depicted in Figure 7, our DuQuant more effectively smooths activations than QuaRot; 2) As demonstrated by the perplexity in Table 7, QuaRot employs GPTQ for their weight quantization method, whereas our DuQuant, with its sophisticated outlier management, attains competitive results using RTN quantization. Additionally, the inclusion of the LWC component within DuQuant significantly enhances the performance of our quantized model. For a more comprehensive comparison, please refer to Appendix F.
@@
Table 8: Quantization runtime on one NVIDIA A100.
Runtime.
@@
To assess the inference speedup delivered by our DuQuant, we adopt the measurement strategy and W4A4 kernel from [2]. We evaluate the layer-wise speedup of LLaMA2-7B and LLaMA2-13B on NVIDIA RTX 3090 GPUs, with results detailed in Table 9. We can observe that during the prefill phase, DuQuant achieves a 2.08×2.08$\times$2.08 × speedup over FP16 for LLaMA2-7B and a 2.34×2.34$\times$2.34 × speedup for LLaMA2-13B, with slight variations across different batch sizes. Additionally, we analyze the computational overhead introduced by rotation and permutation transformations, as illustrated in Figure 9. It can be seen that a single permutation (including one permutation and two rotations) incurs 8.9% on LLaMA2-7B and 9.3% on LLaMA2-13B additional computational costs relative to the W4A4 setup, while setups involving only rotation operations add 1.5% more cost on LLaMA2-7B and 1.8% more cost on LLaMA2-13B. Considering the approximately 2×2$\times$2 × speedup and the impressive performance demonstrated by DuQuant  these additional costs are deemed acceptable. Further permutations do not enhance performance and can negatively impact inference efficiency. More detailed results on inference speedup are available in Appendix E.1.
Table 9: Layer-wise speedup during prefilling phase for 4-bit weight-activation quantization.

### v3

# Linzhan Mou5, Linqi Song4, Zhenan Sun [^2] Corresponding authors. 1,3, Ying Wei22footnotemark: 2 5 1

@@ Abstract
Quantization of large language models (LLMs) faces significant challenges, particularly due to the presence of outlier activations that impede efficient low-bit representation. Traditional approaches predominantly address Normal Outliers, which are activations across all tokens with relatively large magnitudes. However, these methods struggle with smoothing Massive Outliers that display significantly larger values, which leads to significant performance degradation in low-bit quantization. In this paper, we introduce DuQuant, a novel approach that utilizes rotation and permutation transformations to more effectively mitigate both massive and normal outliers. First, DuQuant starts by constructing the rotation matrix, using specific outlier dimensions as prior knowledge, to redistribute outliers to adjacent channels by block-wise rotation. Second, We further employ a zigzag permutation to balance the distribution of outliers across blocks, thereby reducing block-wise variance. A subsequent rotation further smooths the activation landscape, enhancing model performance. DuQuant simplifies the quantization process and excels in managing outliers, outperforming the state-of-the-art baselines across various sizes and types of LLMs on multiple tasks, even with 4-bit weight-activation quantization. Our code is available at https://github.com/Hsu1023/DuQuant.
@@ 1 Introduction
Large language models (LLMs) [52, 7, 51] have demonstrated exceptional performance across a wide range of natural language processing tasks. However, their billions of parameters present considerable deployment challenges on resource-constrained edge devices, particularly in terms of memory usage and inference speed [23, 15, 57]. In response to these challenges, network quantization methods [20, 24] have been extensively explored to minimize memory usage by converting floating-point parameters into low-bit formats [18, 33, 8], and to expedite inference by quantizing both activations and weights for accelerating the matrix multiplication process [65, 35, 76].
@@
As a result, DuQuant offers several clear advantages over QuaRot [2]: (1) DuQuant’s optimal rotation matrix, derived through a greedy search guided by prior knowledge, surpasses QuaRot’s Hadamard rotation in managing outliers; (2) our unique zigzag permutation significantly reduces activation variance across blocks, providing a distinct advantage for handling massive outliers; and (3) by jointly smoothing weights and activations, DuQuant avoids time-consuming GPTQ algorithm in QuaRot. Extensive evaluations demonstrate that our DuQuant approach significantly outperforms existing 4-bit weight-activation quantization baselines across various benchmarks. Notably, DuQuant achieves a 5% improvement in Commonsense QA tasks across all LLaMA model sizes and a 10% increase in zero-shot MMLU benchmarks for the Vicuna-v1.5-13B. Moreover, in practical applications with the LLaMA2-7B model, DuQuant not only accelerates pre-filling phase by up to 2.08$\times$ but also reduces memory usage during decoding phase by 3.50$\times$, with minimal impact on performance: only a 0.61 increase in perplexity and a 2.71% drop in accuracy compared to the FP16 model. These results highlight the effectiveness of DuQuant in enhancing the efficiency and capacity of quantized LLMs.
@@ 2 Motivation
Previous works [13, 74, 33] have highlighted the challenge posed by activation outliers in LLMs for model compression. These outlier features consistently manifest large values across specific feature dimensions and are present in all token sequences [65], which we refer to as Normal Outliers. Recently, a distinct type of outlier [49, 36], termed Massive Outliers, has been observed in LLMs. The primary distinctions between normal and massive outliers are: 1) Normal outliers persist across all token sequences, whereas massive outliers are confined to a limited number of tokens. 2) Massive outliers exhibit significantly larger magnitudes, often surpassing 100 and being approximately 1000 times greater than the median of other activations [49]. In our study, we delve deeper into the impact of these two distinct types of outliers on quantization.
@@
In contrast to previous studies [49, 36] that observe massive outliers at the output of Transformer blocks, we first discover that these extremely large activations exist at the input of the down-projection layer within the FFN module. As depicted in Figure 1, the input of the down-projection layer in the LLaMA2-7B model Layer 1 contains a single activation of significant magnitude (approximately 1400). This activation is isolated to one token and therefore classified as one of massive activations. This phenomenon is consistently observed across different layers and sizes of models, as illustrated in Appendix I.
@@
Although previous studies [65, 48, 40, 1] have proposed various approaches to eliminate outlier features, they still face challenges in effectively managing massive outliers. SmoothQuant [65], for instance, attempts to shift the quantization difficulty from activations to weights by dividing the activation by a per-channel smoothing factor and multiplying it to the weight matrix. Nevertheless, we observe that this transfer at the input of the down-projection layer can cause the weights of the down-projection to display noticeable outliers, as demonstrated in Figure 1 . This issue arises because massive outliers cause the smoothing factor to become significantly large. Moreover, extremely large outliers can lead optimization-based methods to encounter problems with loss explosion. Both OmniQuant [48] and AffineQuant [40] have had to exclude their learnable parameters for the down projection layer due to unstable gradients. Given the poor accuracy observed with 4-bit quantization, QUIK [1] opts to use INT8 quantization for the down projection layer and Atom [76] applies INT8 quantization for 128 outlier channels. Consequently, massive outliers introduce new challenges to the quantization process that existing methods cannot fully address. This observation has motivated us to develop rotation and permutation transformations, which effectively handles both massive and normal outliers and achieves state-of-the-art performance.
@@ 3.2 The proposed DuQuant Method
To address the Normal Outliers issue stated in Section 2, current quantization methods, such as SmoothQuant [65] and OmniQuant [65], usually adopt the smooth technique. Concretely, it involves the utilization of a per-channel smoothing diagonal matrix, denoted as $\mathbf{\Lambda}$, to scale the input activation and weight matrix. The adjustment allows us to rewrite the original linear layer as $\mathbf{Y}=\mathbf{X}\cdot\mathbf{W}=(\mathbf{X}\cdot\mathbf{\Lambda})(\mathbf{\Lambda}^{-1}\cdot\mathbf{W}).$ The diagonal element $\mathbf{\Lambda}_{j}$ within $\mathbf{\Lambda}$ is computed as $\mathbf{\Lambda}_{j}\!=\text{max}(|\mathbf{X}_{j}|)^{\alpha}/\text{max}(|\mathbf{W}_{j}|)^{1-\alpha}$, where $\alpha$ is a hyper-parameter representing the migration strength. However, despite the ability of this smoothing technique to shift the quantization challenge from activations to weights, it still faces difficulties in effectively managing Massive Outliers, as depicted in Figure 1. This challenge stems from the extremely large massive outliers inducing large scaling factors $\mathbf{\Lambda}_{j}$, which in turn introduce new outliers in the weight matrix and result in significant performance declines in 4-bit quantization.
@@
    Let $N$ denote the greedy search steps, then the approximated rotation matrix $\hat{\mathbf{R}}=\mathbf{R}^{1}\mathbf{R}^{2}\cdots\mathbf{R}^{n}$, where $n=\arg\min_{k\in[1:N]}\left(\max_{i,j}|(\mathbf{X}\mathbf{R}^{1}\cdots\mathbf{R}^{k})_{ij}|\right)$. Each $\mathbf{R}^{i}$ is constructed according to Eqn. (2) and the identified feature dimension $d^{(i)}$. Appendix G provides detailed pseudo code.
@@
where the notation $\mathbf{P}$ denotes the orthogonal permutation matrix learned via the zigzag manner, the $\hat{\mathbf{R}}_{(1)}$ and $\hat{\mathbf{R}}_{(2)}$ represent the first and second block-diagonal rotation matrix, respectively.
Remark 1. It is worth noting that the proposed DuQuant method can simultaneously smooth the weight matrix. While the commonly adopted smooth technique is effective, it can cause the weight matrix of the down-projection layer to exhibit pronounced outliers, leading to performance degradation. However, in the proposed DuQuant method, the rotation transformation we designed is applied to not only the activation input but also the weight matrix. As a result, the outliers induced by the smooth technique can be mitigated through our approximated rotation matrix $\hat{\mathbf{R}}$, yielding a smoother, more quantization-friendly weight matrix. Moreover, this approach eliminates the reliance on complex weight quantization techniques, such as GPTQ [18] used in Atom [76] and QuaRot [2].
@@ 4 Experiment
We apply our DuQuant on pre-trained LLMs: LLaMA (7B-65B) [52], LLaMA2 (7B-70B) [53], LLaMA3 (8B, 70B), Mistral, Phi2 and instruction-tuned LLMs: Vicuna-v1.5 (7B-13B) [10]. We evaluate quantized pre-trained LLMs on language generation tasks and commonsense QA tasks. Specifically, we assess the perplexity on WikiText2 [41] and C4 [45] datasets, as well as the zero-shot accuracy on PIQA [6], ARC [12], BoolQ [11], HellaSwag [71], and WinoGrande [46] datasets. Moreover, we evaluate quantized Vicuna models on MMLU [22] and MT-Bench [78] benchmarks, as well as their long-form generative capabilities on LongBench [4].
@@
In line with prior studies [35, 48, 40], we apply per-token activation quantization and per-channel weight quantization. Given that W8A8 quantization has been established as lossless in precision by SmoothQuant [65], our primary evaluation in this paper focuses on 4-bit and 6-bit quantization for weights and activations. As for details, we quantize all intermediate activations, excluding the SoftMax output. Moreover, we have developed two types of quantized models, denoted as DuQuant and DuQuant+LWC . For DuQuant, we employ round-to-nearest quantization, using a clipping ratio of 0.9 for activations and 0.8 for weights. To improve weight matrix quantization, DuQuant+LWC integrates the learnable weight clipping (LWC) technique from OmniQuant. Concretely, LWC adjusts weights by training parameters $\gamma,\beta\in[0,1]$ to compute step size $\Delta=\frac{\gamma\max(\mathbf{X})-\beta\min(\mathbf{X})}{2^{b}-1}$ in Eqn. (1). Notably, the smoothing diagonal matrix and the learned weight clipping factor can be integrated into the quantized weights, introducing no additional computational or memory costs. More details and hyperparameters are left in Appendix C.
@@
Table 2: Zero-shot QA ($\uparrow$) results of LLaMA1 models under 4-bit weight-activation quantization. The results for LLaMA2 models and W6A6 quantization can be found in Table D1  D9, and  D10.
@@ 4.1 Main Results
We conduct a comprehensive comparison of our DuQuant with several SOTA baselines on LLaMA1 and LLaMA2 models. Results for W4A4 quantization are presented in this Section, while results for W6A6 quantization are provided in Appendix D. Table 1 indicates that our DuQuant quantized models notably outperform other baselines on both the WikiText2 and C4 datasets. Notably, LWC technique further enhances model capacity, with our DuQuant+LWC achieving comparable performance with FP16 models. Table 2 and Table D1 showcase the zero-shot accuracy of W4A4 quantization on Commonsense QA tasks, where DuQuant significantly improves the average accuracy. Our method surpasses QLLM by +9%, and Atom by +5% for all model sizes. These results demonstrate the superiority of our rotation and permutation transformation, which establishes new SOTA performance by effectively eliminating outlier features.
@@
Table 4: Long-context generation results for 4-bit Vicuna models on the LongBench benchmark.
Vicuna Setting Qasper QMSum MultiNews TREC TriviaQA SAMSum DuReader RepoBench-P Avg
Vicuna-v1.5-7B W4A4 FP16 23.27 21.07 26.91 66.00 82.59 41.06 25.53 48.23 41.83
SmoothQuant 4.11 2.00 6.05 15.00 1.62 1.55 4.24 25.92 7.56
OmniQuant 1.62 3.93 2.64 1.00 0.81 0.61 1.87 14.97 3.43
Atom 17.97 20.24 24.60 58.00 67.20 37.94 19.41 29.34 34.34
DuQuant 19.98 21.15 25.85 64.00 78.91 42.24 23.15 47.66 40.37
Vicuna-v1.5-13B W4A4 FP16 24.41 21.24 26.53 68.00 86.81 41.97 27.57 43.08 42.45
SmoothQuant 2.18 2.95 3.54 1.50 1.83 0.35 6.71 11.57 3.83
OmniQuant 0.68 1.78 2.83 9.00 1.13 0.45 13.83 8.46 4.77
Atom 17.67 20.23 23.39 59.00 80.75 38.72 21.79 37.31 37.36
DuQuant 18.93 20.72 26.59 66.50 83.04 42.67 26.02 38.09 40.32
@@
Evaluation of Long-context Generation.
To further evaluate the long-text generative capabilities, we follow  [38, 34, 56] and conduct a comprehensive comparison of DuQuant against state-of-the-art baselines on the LongBench [4], which includes a variety of generative tasks to provide a broader evaluation. We set the maximum sequence length to 3500 for Vicuna models, with results presented in Table 4. DuQuant achieves performance comparable to FP16 models, demonstrating the effectiveness of our dual transformations. More detailed results on different subtasks are listed in Table D3,  D4.
@@ 4.2 Ablation Study
We ablate four distinct operations within DuQuant: 1) only the smoothing technology like SmoothQuant; 2) one rotation following the smoothing operation; 3) a sequence of rotation, permutation, and another rotation without smoothing; and k4) full DuQuant approach. Table 6 shows that the smoothing operation plays a basic role in our DuQuant by shifting activation outliers to weight. The initial rotation significantly enhances model performance, yielding competitive PPL results. Finally, permutation combined with a second rotation further enhances the quantized model.
@@
In light of the recent introduction of Hadamard rotations by QuaRot[2] to eliminate outlier features, we have undertaken a detailed analysis to highlight the key differences between our DuQuant and QuaRot. To ensure a balanced evaluation, we have re-implemented QuaRot in accordance with our quantization settings. The results demonstrate that 1) the rotation matrix constructed by DuQuant outperforms QuaRot’s approach of simply selecting a randomly initialized Hadamard matrix. As depicted in Figure 11, our DuQuant more effectively smooths activations than QuaRot. This is attributed to the prior knowledge utilized by DuQuant to accurately target the outliers; 2) As demonstrated by the perplexity in Table 8, QuaRot employs GPTQ for their weight quantization method, whereas our DuQuant, with its sophisticated outlier management, attains competitive results using RTN quantization. The superiority proves the effectiveness of our zigzag permutation to enhance capacity. For a more comprehensive comparison, please refer to Appendix F.
@@
Permutation Frequency.
We conduct ablations on the rotation and permutation frequencies in DuQuant. As shown in Figure 9, “Perm 1” (two rotations with one permutation) achieves stronger performance compared with “Perm 0” (no permutation), while incurring an additional 8.9% computational cost on LLaMA2-7B and 9.3% on LLaMA2-13B compared to the W4A4 setup. Considering the approximately 2×2$\times$2 × speedup and the impressive performance, these additional costs are deemed acceptable. Further permutations, like “Perm 2,” do not improve performance and reduce inference efficiency. Consequently, “Perm 1” strikes the best balance between perplexity and inference speed, making it the optimal configuration for DuQuant.
@@
To assess the inference speedup delivered by our DuQuant, we adopt the measurement strategy and W4A4 kernel from [2]. We evaluate the layer-wise speedup of LLaMA2 models on one NVIDIA 3090 GPU, with results detailed in Table 9 and 11. We set the pre-filling sequence length at 2048 and decode for 128 steps. In the pre-filling stage, DuQuant achieves a 2.08×2.08$\times$2.08 × speedup over FP16 for LLaMA2-7B and a 2.34×2.34$\times$2.34 × speedup for LLaMA2-13B, with slight variations across different batch sizes. In the decoding stage, batching the token generation phase yields high throughput without any downside [44]. Consequently, we enlarge the batch size to 64 and the results for LLaMA2-7B in Table 11 prove DuQuant achieves speedup comparable to QuaRot. More detailed analyses and end-to-end speedup are available in Appendix E.1.
Table 9: Layer-wise speedup during pre-filling stage for 4-bit weight-activation quantization.
@@
Memory Consumption.
We measure the peak memory usage of DuQuant with the W4A4 kernel on LLaMA2-7B using a single NVIDIA 3090 GPU. We process 2048 tokens for pre-filing and run 128 decoding steps, with the results listed in Table 11. In the pre-filling stage, DuQuant, SmoothQuant, and QuaRot achieve up to 3.2$\times$ memory reduction, while QLLM performs worse. In the decoding stage, DuQuant maintains strong memory efficiency, with superior performance.
Table 10: Decoding stage.
INT4, BS=64 Speedup
FP16 -
SmoothQuant 1.508$\times$
QLLM OOM
QuaRot 1.442$\times$
DuQuant 1.321$\times$
Table 11: Peak memory usage with a batch size of 1.
LLaMA2-7B Pre-filling (GB) Saving Decoding (GB) Saving
FP16 15.282 - 13.638 -
SmoothQuant 4.782 3.196$\times$ 3.890 3.506$\times$
QLLM 5.349 2.857$\times$ 3.894 3.502$\times$
QuaRot 4.784 3.194$\times$ 3.891 3.505$\times$
DuQuant 4.786 3.193$\times$ 3.893 3.503$\times$
Table 12: Quantization runtime on one NVIDIA A100.
Runtime.